Monday, November 11, 2002

working on the upgrade at work, logging the progress, I was reminded that I need to blog!! Here's that log...
-----Original Message-----
From: Jan
Sent: Sunday, November 10, 2002 9:30 PM
To: Techteam; Peggy Teal
Subject: CIRC Hardware BLOG 11/10/02

x Tedious detail ALERT! x
If you are sensitive, delete or file now!

Here's how it went today, Sunday, 11/10/02...

12:15 I arrive to do the daily backup manually.

12:35 pounded out cron, locking the databases, timing the backup...

2:10 the directories are backed up to the tape, 1:40 elapsed time

3:05 the verify is done and tape ejected, total for all steps about 2-1/2

3:40 Norm gets here. First, he installed all the new memory

4:30 problem with ALL the hard drives
Part #76h2697 for a 4g drive on the top row w/white face plate (in the #1

There are 4 on middle row v/ black face plate. On the middle row, there are
6 slots left to rgt (X= drive in slot, O= open)--
1 2 3 4 5 6

There is no boardfor drives on the bottom row.

At this time, we went with staying with all the old drives. NO NEW ONES IN
THE SERVER. 2 of the new drives face the right way. We can get Summitt to
send a back set (or 4?) that will fit overnight tomorrow. We can install,
turn key, open front and drives slide right in.

Need to hear from Summit and PBS about what parts are right and make a new
plan. I am hoping to get this poked in there and back on schedule no later
than the end of the day Wednesday.

5:10 memory problem. F50 will not come back up. Norm calls support. They
confirm that new memory board should have gone in original position and new
one (because it is full w/16)and the old one (it has 12 slots full only)
should qo into next position. Norm tries to switch. System cannot even
tell us which of the chips may be bad,reseatin chips.. Try again.

5:37 we have a login.

5:47 Running diag, Norm runs into trouble with F7 key. Could be emulation.
But we switch keybds w/ no effect. Calls IBM for help.

6:20 I asked Norm if we could telnet from Adv, since it is ibm3151
emulation, we did, F7 -- this worked (I don't know if this happened to us

6:25 We have two bad chips, Norm is taking these out and moving them all
back up in sequence.

8:25 At last the diag errors were gone, after Norm reseat the chips, moved
'em around and restarted the system about 8 times.

8:35 I have the system back, checked lpstat, unlocked db and checked
application and spool

8:38 calling PBS

8:55 Gary dialed in, will apply the multivolume to the databases that need
it, will be done before 10 and will unlock the databases for the users.
Gary verifies that we now have 3GB memory, yay!

9:20 Norm called back, wanted work order and contract #!

9:30 I am going home, remind me that the cron needs to be set back so it
will run tomorrow night.

No comments: